OpenStudy (howard-wolowitz):

Explanation needed here:

OpenStudy (jtvatsim):

Did that other weird one end up being 5 as the correct answer? @Howard-Wolowitz

OpenStudy (howard-wolowitz):

havent graded it yet

OpenStudy (howard-wolowitz):

last one

OpenStudy (jtvatsim):

k... doesn't help then :)

OpenStudy (howard-wolowitz):

what you thinking loser

OpenStudy (howard-wolowitz):

isnt the formula L=A or A = something

OpenStudy (loser66):

I have no idea!! to me cosecant function has NO amplitude

OpenStudy (howard-wolowitz):

thats what another guy said

OpenStudy (loser66):

This is what I got to make sure what I thought http://www.cathedralcatholic.org/faculty-resources/82/63/Ch%204%20Notes%20Part%202.pdf

OpenStudy (loser66):

the last paragraph

OpenStudy (jdoe0001):

yeap, I agree with Loser66, quite unsure on what it's being asked as well

OpenStudy (jdoe0001):

the function has a cyclical range, and thus no height per se, not to mention has cyclical asymptotes as well

OpenStudy (jtvatsim):

If we interpret the phrase "height" naively, we would probably say that the height is 2 (that is the amplitude in front of the trig function). Frankly, none of the choices seems to make sense, but that is as good as I would do myself.

OpenStudy (xapproachesinfinity):

can't even see the picutre

OpenStudy (jtvatsim):

@xapproachesinfinity click on the picture to magnify.

OpenStudy (howard-wolowitz):

well i dont know dude

OpenStudy (xapproachesinfinity):

does not appear even if i magnify it

OpenStudy (jtvatsim):

QUESTION: "Identify the height for the function: \[f(t) = 2\csc\left(4t - \frac{\pi}{5} \right) + 1\] Choices: A) 2pi/5 B) 8 C) 2 D) 3"

OpenStudy (howard-wolowitz):

thanks!

OpenStudy (xapproachesinfinity):

now i can see it the height of \[\csc x=1\] so \[\csc(4t-\frac{\pi}{5})=1 \\ 2\csc (4t-\pi/5)=2\] \[2\csc(4t-\pi/5)+1=3\] \[y=3\] that would be the height

OpenStudy (howard-wolowitz):

hmmm

OpenStudy (xapproachesinfinity):

actually i do agree with loser csc has no amplitude lol

OpenStudy (xapproachesinfinity):

perhaps they meant what is the local minima or something

OpenStudy (howard-wolowitz):

well thanks for trying

OpenStudy (xapproachesinfinity):

i had the feeling of using local minima that happens when cscx=1

OpenStudy (xapproachesinfinity):

what course is this ?

OpenStudy (howard-wolowitz):

advanced algebra

OpenStudy (xapproachesinfinity):

your course is wrongly interpreting the concepts of height i saw that in your previous questions as well

OpenStudy (xapproachesinfinity):

@jim_thompson5910

OpenStudy (xapproachesinfinity):

as far as i know i agree with loser and jdoe

OpenStudy (howard-wolowitz):

@zepdrix

OpenStudy (howard-wolowitz):

i gotta help Ur help @zepdrix

zepdrix (zepdrix):

Height is usually measured from the x-axis, so I would imagine either 2 or 3 is the correct answer. I would lean towards 3, but again, I agree with everyone else here, the wording is very poor.

OpenStudy (howard-wolowitz):

thank you for helping